# MMS-LLaMA: Efficient LLM-based Audio-Visual Speech Recognition with Minimal Multimodal Speech Tokens

### v1

@@ Abstract
Audio-Visual Speech Recognition (AVSR) achieves robust speech recognition in noisy environments by combining auditory and visual information. However, recent Large Language Model (LLM) based AVSR systems incur high computational costs due to the high temporal resolution of audio-visual speech processed by LLMs. In this work, we introduce an efficient multimodal speech LLM framework that minimizes token length while preserving essential linguistic content. Our approach employs an early av-fusion module for streamlined feature integration, an audio-visual speech Q-Former that dynamically allocates tokens based on input duration, and a refined query allocation strategy with a speech rate predictor to adjust token allocation according to speaking speed of each audio sample. Extensive experiments on the LRS3 dataset show that our method achieves state-of-the-art performance with a WER of 0.74% while using only 3.5 tokens per second. Moreover, our approach not only reduces token usage by 86% compared to the previous multimodal speech LLM framework, but also improves computational efficiency by reducing FLOPs by 35.7%. The code and models are available https://github.com/JeongHun0716/MMS-LLaMA.
@@ 1 Introduction
In human communication, watching lip movements and listening to sounds are essential for understanding speech. These multimodal cues, which combine visual and auditory information, enable people to communicate effectively in bustling cafes, crowded streets, and noisy factories. Thanks to these practical advantages and significant advances in deep learning, Audio-Visual Speech Recognition (AVSR) technology has made remarkable progress through numerous research efforts Shi et al. (2022a); Ma et al. (2023); Afouras et al. (2018a); Ma et al. (2021); Serdyuk et al. (2022); Cappellazzo et al. (2024). Now, it is easy to find an AVSR model that can accurately predict what you have said, even in noisy environments.
@@
Rapid progress in audio-visual speech recognition has been driven by large-scale audio-visual datasets Afouras et al. (2018b, a), advanced neural architectures Elman (1990); Vaswani (2017); Gulati et al. (2020), multimodal learning strategies (e.g., self-supervised learning Shi et al. (2022a) and knowledge distillation from pre-trained ASR Ma et al. (2023)), carefully designed training methods Ma et al. (2022); Hong et al. (2023), and the integration of Large Language Models (LLMs) for sentence prediction Cappellazzo et al. (2024). Among these, multimodal speech LLM frameworks stand out by integrating LLMs for enhanced context modeling. However, despite these gain, they still suffer from high computational costs. This is largely due to multimodal speech tokens having a higher temporal resolution compared to text tokens, which forces the self-attention mechanism in each LLM layer to process many more tokens, thereby significantly increasing the computational burden.
@@
To address these issues, we propose an efficient multimodal speech LLM framework that preserves essential linguistic information while significantly reducing the length of multimodal speech tokens. We refer to this framework as Multimodal Speech-LLaMA (MMS-LLaMA), which is specifically designed for AVSR tasks. We construct this framework with three primary components: 1) early av-fusion module shifts the fusion process to an earlier stage, prior to inputting multimodal speech tokens into the LLM. 2) audio-visual speech Q-former (AV Q-Former) is designed to dynamically allocate the number of learnable queries according to the duration of audio-visual input, where queries are transformed into multimodal speech tokens. 3) Going one step further, to more effectively adjust the number of multimodal tokens, we propose a speech rate predictor. It accounts for differences in speaking speed across each audio sample, enabling the assignment of more tokens to faster speech.
@@
Reflecting the growing trend toward efficient multimodal models and leveraging insights from efficient Large Multimodal Models (LMMs) that utilize minimal vision tokens during pre-fusion Zhang et al. (2025), we incorporate their effective pre-fusion approach into our multimodal speech LLM framework. Concurrently, we explore and integrate existing audio-visual fusion strategies into our early-AV fusion module, which reduces the number of fused sequences by half.
@@
Through extensive experiments, we demonstrate that using only 3.5 multimodal speech tokens per second can effectively preserve language content while maintaining performance. Moreover, our approach achieves state-of-the-art performance with a Word Error Rate (WER) of 0.74% on the LRS3 dataset. Compared to the previous LLM-based multimodal speech framework Cappellazzo et al. (2024) that uses 25 tokens per second, our method reduces token usage by 86%. Additionally, our approach improves computational efficiency, reducing FLOPs by 35.7%.
@@ 2.1 Audio-Visual Speech Recognition
Audio-based Automatic Speech Recognition (ASR) is a well-studied and one of the most frequently used technology Amodei et al. (2016); Kim et al. (2017); Prabhavalkar et al. (2023). However, as it primarily depends on acoustic inputs, its performance naturally degraded when the input audio is perturbed with other background noises (e.g., babble noise) Varga and Steeneken (1993). Visual Speech Recognition (VSR) aims to employ visual inputs only in speech recognition, using the fact that the visual inputs are not affected by the acoustic noises. A diverse range of literature has been explored on VSR techniques to avoid the impact of background noise on speech recognition performance Petridis et al. (2017, 2018b); Martinez et al. (2020); Ma et al. (2021, 2022, 2023); Kim et al. (2022, 2023, 2024a, 2024b); Yeo et al. (2024b, c, d).
Audio-Visual Speech Recognition (AVSR) integrates both audio and visual modalities as input, combining the strengths of ASR and VSR techniques to enhance the robustness and performance of speech recognition systems, particularly in noisy environments. Early AVSR models Noda et al. (2015); Huang and Kingsbury (2013); Mroueh et al. (2015); Stewart et al. (2013) established the foundation for multimodal speech recognition, demonstrating the effectiveness of fusing visual features such as lip movements with audio signals. Advancements in AVSR Petridis et al. (2018a); Shi et al. (2022b); Hong et al. (2022, 2023) have been fueled by improvements in both data availability Chung et al. (2017); Afouras et al. (2018b) and model architectures Vaswani (2017); Gulati et al. (2020). Self-supervised learning has also played a crucial role in advancing AVSR Shi et al. (2022a); Lian et al. (2023); Haliassos et al. (2024b, a) to further enhance performance.
@@ 2.2 Speech Large Language Model
Early advancements in speech recognition began with monolingual English ASR models Hsu et al. (2021); Radford et al. (2023). Building on these monolingual models, multilingual speech recognition systems Fathullah et al. (2024); Rubenstein et al. (2023) have significantly improved performance, especially for low-resource languages, by leveraging the multilingual capabilities of LLMs. Building on the contextual understanding capabilities of LLMs, current Speech LLMs have significantly greatly improved the accuracy, robustness to noise, and adaptability to diverse accents and dialects. Beyond speech recognition, Speech LLMs have also expanded to support multitask learning, enabling a single model to perform a wide range of speech-related tasks Chu et al. (2023); Tang et al. (2023). Qwen-Audio Chu et al. (2023) and SALMONN Tang et al. (2023) scale audio-language pre-training to cover various speech-related tasks and diverse audio inputs. Despite these significant progresses, the exploration of extending LLMs’ capabilities to the audio-visual speech domain remains limited.
However, most current speech LLMs primarily focus on audio-based tasks and have not explored the effectiveness of adopting utilization of speech rate. In contrast, our work incorporates a speech rate predictor that dynamically allocates resources based on speaking speed of each audio, enabling more efficient and robust processing for audio-visual speech inputs.
@@ 3.3 Speech Rate Predictor
  Method Audio Encoder Visual Encoder Decoder Training Data(hrs) WER $\downarrow$
@@
Whisper AV-HuBERT LLaMA 3.2 3B 1759 1.9 0.74
Table 1: Comparisons with state-of-the-art methods on the LRS3 dataset. We report each method’s architecture (audio encoder, visual encoder, decoder), the amount of training data, token throughput, and WER under both clean and noisy conditions. The clean condition is evaluated on the original test set, while the noisy condition is evaluated on a test set with babble noise added at 0-SNR
@@ 4.1 Dataset
Lip Reading Sentences 3 (LRS3) as detailed in Afouras et al. (2018b), is a widely used dataset designed for audio-visual speech recognition. It includes 433 hours of audio-visual data sourced from TED and TEDx talks, accompanied by human-annotated text transcriptions.
VoxCeleb2 as detailed in Chung et al. (2018), is a dataset designed for speaker recognition. It consists of 2,442 hours of multilingual audio-visual data. Following Shi et al. (2022a), we utilize only the English portion of this dataset, which amounts to 1,326 hours. Moreover, we also use the Whisper ASR model to generate pseudo text transcriptions, which we combine with the LRS3 dataset for training our model. This combined dataset amounts to 1,756 hours.
@@ 5.1 Comparison with the state-of-the-art methods
Traditional approaches, such as CM-seq2seq Ma et al. (2021), ViT3D-CM Serdyuk et al. (2022), and auto-avsr Ma et al. (2023), commonly employ Conformer- or Transformer-based backbones and rely on large-scale datasets to improve AVSR performances. Notably, LP Conformer Chang et al. (2024) achieved a 0.9% WER using 100K hours of training data. More recent models, including Whisper-Flamingo Rouditchenko et al. (2024) and LLaMA-AVSR Cappellazzo et al. (2024), leverage large-scale pretrained models (e.g., Whisper and LLMs) and have demonstrated superior performance with WERs of 0.77% and 0.76%, respectively. Our MMS-LLaMA surpasses previous state-of-the-art methods, achieving a WER of 0.74% when trained on 1,759 hours of data. When trained on only 433 hours of data, MMS-LLaMA obtains a WER of 0.92%, outperforming both Whisper-Flamingo and LLaMA-AVSR. Under noisy conditions, our model also achieves state-of-the-art performance, with a WER of 1.9%. Notably, our model significantly improves performance in noisy conditions, addressing a major weakness of previous multimodal speech LLM framework that showed a substantial gap between noisy and clean conditions. Please note that the proposed MMS-LLaMA not only achieves the superior performances but also effectively reduces the number of tokens with the proposed AV Q-Former.
Method # MMS Tokens per second GPU Memory Usage(GB) FLOPs (T) WER(%)
@@ 5.2.1 Validation of the Effectiveness of Each Component via Sequential Integration
Types of AV-Fusion FLOPs(T) WER $\downarrow$
@@
Similarly, we incorporate the speech rate predictor at the same query frequency settings, to validate its effectiveness. Specifically, at a query frequency of 1, the FLOPs and GPU memory usage increases by just 0.1T and 0.1GB, leading to a WER improvement to 1.61%. At frequencies of 2 and 3, the FLOPs slightly incrase at 1.40T and 1.44T, respectively, and GPU memory usage is nearly the same, while the WER improves further to 0.97% and 0.92%. Furthermore, the analysis shows that token usage increases by only 0.2, 0.4, and 0.7 tokens per second for query frequencies of 1, 2, and 3, respectively. These minor increases do not significantly affect FLOPs, yet still yield WER improvements of 0.38%, 0.13%, and 0.03% under the same settings.
@@ 5.2.2 Evaluation of Different Audio-Visual Fusion Strategies
The results, presented in Table 3, indicate that all three fusion techniques exhibit similar FLOPs. Under noisy conditions, concatenation achieves the best WER of 2.77%, while multimodal attention performs best on clean speech with 0.87% WER, followed by concatenation of 0.92% WER. Because the performance gap in noisy settings is more significant than in clean conditions, we adopt concatenation as our primary fusion strategy in the other experiments.
@@ 5.2.3 Impact of LLM Model Size
As shown in Table 4, the LLaMA3.2 3B model achieved the best WER of 0.92% under clean conditions, outperforming both the LLaMA3.2 1B and LLaMA3.1 8B models. However, under noisy conditions, the LLaMA3.1 8B model achieved the best WER of 2.61%. These results suggest that larger model might perform better under the challenging environment. As expected, the larger model incurs higher GPU memory usage and computational costs. Given its comparable performance to the 8B model, we report results using the LLaMA3.2-3B model for other experiments due to its significantly lower computational requirements.

### v2

@@ Abstract
Audio-Visual Speech Recognition (AVSR) achieves robust speech recognition in noisy environments by combining auditory and visual information. However, recent Large Language Model (LLM) based AVSR systems incur high computational costs due to the high temporal resolution of audio-visual speech processed by LLMs. In this work, we introduce an efficient multimodal speech LLM framework that minimizes token length while preserving essential linguistic content. Our approach employs an early AV-fusion module for streamlined feature integration, an audio-visual speech Q-Former that dynamically allocates tokens based on input duration, and a refined query allocation strategy with a speech rate predictor to adjust token allocation according to speaking speed of each audio sample. Extensive experiments on the LRS3 dataset show that our method achieves state-of-the-art performance with a WER of 0.72% while using only 3.5 tokens per second. Moreover, our approach not only reduces token usage by 86% compared to the previous multimodal speech LLM framework, but also improves computational efficiency by reducing FLOPs by 35.7%. The code and models are available https://github.com/JeongHun0716/MMS-LLaMA.
@@ 1 Introduction
In human communication, watching lip movements and listening to sounds are essential for understanding speech. These multimodal cues, which combine visual and auditory information, enable people to communicate effectively in bustling cafes, crowded streets, and noisy factories. Thanks to these practical advantages and significant advances in deep learning, Audio-Visual Speech Recognition (AVSR) technology has made remarkable progress through numerous research efforts Shi et al. (2022a); Hu et al. (2022); Ma et al. (2023); Afouras et al. (2018a); Ma et al. (2021); Serdyuk et al. (2022); Fu et al. (2024); Cappellazzo et al. (2024). Now, it is easy to find an AVSR model that can accurately predict what you have said, even in noisy environments.
@@
Rapid progress has been driven by large-scale audio-visual datasets Afouras et al. (2018b, a), advanced neural architectures Elman (1990); Vaswani (2017); Gulati et al. (2020), improved multimodal learning strategies, including self-supervised learning Shi et al. (2022a) and knowledge distillation using a pre-trained Automatic Speech Recognition (ASR) model Ma et al. (2023); carefully designed training methods Ma et al. (2022); Hong et al. (2023); and the utilization of Large Language Models’ (LLMs) language understanding capabilities as sentence predictors Cappellazzo et al. (2024). Among these, multimodal speech LLM frameworks stand out by integrating LLMs for enhanced context modeling. However, despite these gain, they still suffer from high computational costs. This is largely due to multimodal speech tokens having a higher temporal resolution compared to text tokens, which forces the self-attention mechanism in each LLM layer to process many more tokens, thereby significantly increasing the computational burden.
@@
To address these issues, we aim to develop an efficient MultiModal Speech LLM framework (MMS-LLaMA) that minimizes the lenght of multimodal speech tokens while preserving their linguistic content. To achieve this, we construct this framework with three primary components: 1) early AV-fusion module shifts the fusion process to an earlier stage, prior to inputting multimodal speech tokens into the LLM. 2) Audio-Visual speech Q-former (AV Q-Former) is designed to dynamically allocate the number of learnable queries according to the duration of audio-visual input, where queries are transformed into multimodal speech tokens. 3) Going one step further, to more effectively adjust the number of multimodal tokens, we propose a speech rate predictor. It accounts for differences in speaking speed across each audio sample, enabling the assignment of more tokens to faster speech.
@@
Reflecting the growing trend toward efficient multimodal models and leveraging insights from efficient multimodal LLMs that utilize minimal vision tokens during pre-fusion Zhang et al. (2025), we incorporate their effective pre-fusion approach into our multimodal speech LLM framework. Concurrently, we explore and integrate existing audio-visual speech fusion strategies into our early-AV fusion module, which reduces the number of fused sequences by half.
@@
Through extensive experiments, we demonstrate that using only 3.5 multimodal speech tokens per second can effectively preserve language content while maintaining performance. Moreover, our approach achieves state-of-the-art performance with a Word Error Rate (WER) of 0.72% on the LRS3 Afouras et al. (2018b) dataset. Compared to the previous LLM-based AVSR model Cappellazzo et al. (2024) that uses 25 tokens per second, our method reduces token usage by 86%. Additionally, our approach improves computational efficiency, reducing FLOPs by 35.7%.
@@ 2.1 Audio-Visual Speech Recognition
Audio-based automatic speech recognition is a well-studied and one of the most frequently used technology Amodei et al. (2016); Kim et al. (2017); Prabhavalkar et al. (2023). However, as it primarily depends on acoustic inputs, its performance naturally degraded when the input audio is perturbed with other background noises (e.g., babble noise) Varga and Steeneken (1993). Visual Speech Recognition (VSR) aims to employ visual inputs only in speech recognition, using the fact that the visual inputs are not affected by the acoustic noises. A diverse range of literature has been explored on VSR techniques to avoid the impact of background noise on speech recognition performance Petridis et al. (2017, 2018b); Martinez et al. (2020); Ma et al. (2021, 2022, 2023); Kim et al. (2022); Cheng et al. (2023b); Kim et al. (2023, 2024a, 2024b); Yeo et al. (2024b, c, d).
Audio-visual speech recognition integrates both audio and visual modalities as input, combining the strengths of ASR and VSR techniques to enhance the robustness and performance of speech recognition systems, particularly in noisy environments. Early AVSR models Noda et al. (2015); Huang and Kingsbury (2013); Mroueh et al. (2015); Stewart et al. (2013) established the foundation for multimodal speech recognition, demonstrating the effectiveness of fusing visual features such as lip movements with audio signals. Advancements in AVSR Petridis et al. (2018a); Shi et al. (2022b); Hong et al. (2022, 2023) have been fueled by improvements in both data availability Chung et al. (2017); Afouras et al. (2018b) and model architectures Vaswani (2017); Gulati et al. (2020). Self-supervised learning has also played a crucial role in advancing AVSR Shi et al. (2022a); Lian et al. (2023); Haliassos et al. (2024b, a) to further enhance performance.
@@ 2.2 Speech Large Language Model
Advancements in speech recognition have been achieved through self-supervised learning Hsu et al. (2021) and the development of large-scale ASR datasets Radford et al. (2023). Building on these models, more recent speech recognition systems Fathullah et al. (2024); Rubenstein et al. (2023) have significantly improved performance, especially for low-resource languages, by leveraging the multilingual capabilities of LLMs. Additionally, by utilizing LLM’s capabilities, error correction methods Chen et al. (2023, 2024); Radhakrishnan et al. (2023) have further enhanced the reliability and accuracy of ASR outputs. Beyond speech recognition, Speech LLMs have also expanded to support multitask learning, enabling a single model to perform a wide range of speech-related tasks Chu et al. (2023); Tang et al. (2023). Qwen-Audio Chu et al. (2023) and SALMONN Tang et al. (2023) scale audio-language pre-training to cover various speech-related tasks and diverse audio inputs.
Despite this significant progress, the exploration of extending LLMs’ capabilities to the audio-visual speech domain remains limited. Moreover, most current speech LLMs primarily focus on audio-based tasks and have under-explored the effectiveness of adopting the utilization of speech rate. In contrast, our work incorporates a speech rate predictor that dynamically allocates resources based on the speaking speed of each audio, enabling more efficient and robust processing for audio-visual speech inputs.
@@ 3.3 Speech Rate Predictor
  Method Audio Encoder Visual Encoder Decoder Training Data(h) WER(%$\downarrow$
@@
Whisper AV-HuBERT LLaMA 3.2 3B 1759 1.9 0.72
Table 1: Comparisons with state-of-the-art methods on the LRS3 dataset. We report each method’s architecture (audio encoder, visual encoder, decoder), the amount of training data, and WER under both clean and noisy conditions. The clean condition is evaluated on the original test set, while the noisy condition is evaluated on a test set with babble noise added at 0-SNR
@@ 4.1 Dataset
Lip Reading Sentences 3 (LRS3), as detailed in Afouras et al. (2018b), is a widely used dataset designed for AVSR. It includes 433 hours of audio-visual data for training and 1 hour for evaluation. The videos are sourced from TED and TEDx talks and are accompanied by human-annotated text transcriptions.
VoxCeleb2 as detailed in Chung et al. (2018), is a dataset designed for speaker recognition. It consists of 2,442 hours of multilingual audio-visual data. Following Shi et al. (2022a), we utilize only the English portion of this dataset, which amounts to 1,326 hours. Moreover, we also use the Whisper ASR model to generate pseudo text transcriptions, which we combine with the LRS3 dataset for training our model. This combined dataset amounts to 1,759 hours.
@@ 5.1 Comparison with the state-of-the-art methods
Traditional approaches, such as CM-seq2seq Ma et al. (2021), ViT3D-CM Serdyuk et al. (2022), and auto-avsr Ma et al. (2023), commonly employ Conformer- or Transformer-based backbones and rely on large-scale datasets to improve AVSR performances. Notably, LP Conformer Chang et al. (2024) achieved a 0.9% WER using 100K hours of training data. More recent models, including Whisper-Flamingo Rouditchenko et al. (2024) and LLaMA-AVSR Cappellazzo et al. (2024), leverage large-scale pretrained models (e.g., Whisper and LLMs) and have demonstrated superior performance with WERs of 0.77% and 0.76%, respectively. Our MMS-LLaMA surpasses previous state-of-the-art methods, achieving a WER of 0.72% when trained on 1,759 hours of data. When trained on only 433 hours of data, MMS-LLaMA obtains a WER of 0.9%, outperforming both Whisper-Flamingo and LLaMA-AVSR. Under noisy conditions, our model also achieves state-of-the-art performance, with a WER of 1.9%. Notably, our model significantly improves performance in noisy conditions, addressing a major weakness of previous multimodal speech LLM framework that showed a substantial gap between noisy and clean conditions. Please note that the proposed MMS-LLaMA not only achieves the superior performances but also effectively reduces the number of tokens with the proposed AV Q-Former.
Method # MMS Tokens per second GPU Memory Usage(GB) FLOPs (T) WER(%$\downarrow$
@@ 5.2.1 Validation of the Effectiveness of Each Component via Sequential Integration
Types of AV-Fusion FLOPs(T) WER(%$\downarrow$
@@
Similarly, we incorporate the speech rate predictor at the same query frequency settings, to validate its effectiveness. Specifically, at a query frequency of 1, the FLOPs and GPU memory usage increases by just 0.1T and 0.1GB, leading to a WER improvement to 1.61%. At frequencies of 2 and 3, the FLOPs slightly incrase at 1.40T and 1.44T, respectively, and GPU memory usage is nearly the same, while the WER improves further to 0.97% and 0.9%. Furthermore, the analysis shows that token usage increases by only 0.2, 0.4, and 0.7 tokens per second for query frequencies of 1, 2, and 3, respectively. These minor increases do not significantly affect FLOPs, yet still yield WER improvements of 0.38%, 0.13%, and 0.05% under the same settings.
@@ 5.2.2 Evaluation of Different Audio-Visual Fusion Strategies
The results, presented in Table 3, indicate that all three fusion techniques exhibit similar FLOPs. Under noisy conditions, concatenation achieves the best WER of 2.4%, while multimodal attention performs best on clean speech with 0.87% WER, followed by concatenation of 0.9% WER. Because the performance gap in noisy settings is more significant than in clean conditions, we adopt concatenation as our primary fusion strategy in the other experiments.
@@ 5.2.3 Impact of LLM Model Size
As shown in Table 4, the LLaMA3.2 3B model achieved the best WER of 0.9% under clean conditions, outperforming both the LLaMA3.2 1B and LLaMA3.1 8B models. Under noisy conditions, it also achieved the lowest WER of 2.4%. As expected, increasing the model size leads to higher GPU memory usage and computational costs. Given that the 3B model achieves the best performance with a balanced trade-off in computational cost, we use LLaMA3.2-3B for other experiments.
@@ 5.2.4 Performance Across SNR Levels and Query Frequencies
5.3 Audio-Unavailable Scenarios
5.3.1 Speech Rate Predictor Ablation
In real-world scenarios, audio signals are often distorted due to various factors. To mitigate this issue, we have trained a visual speech rate predictor that relies on lip movements.
To confirm the effectiveness of the visual speech predictor, we have replaced the audio-based speech predictor with the visual speech predictor in our AVSR framework and evaluated its performance on a pre-trained model. The results, shown in Table 6, indicate that the visual speech rate predictor achieves a WER of 0.75%, which is nearly the same as the 0.72% obtained by the audio-based predictor.
Method WER (%)
w/ Speech rate predictor 0.72
w/ Visual speech rate predictor 0.75
Table 6: Comparison between audio-based and visual speech rate predictors.
5.3.2 VSR performance
To evaluate the effectiveness of the proposed method on audio-unavailable scenarios, we have conducted an additional experiment in which the proposed method was trained solely on the visual modality. In this experiment, we have utilized a visual speech rate predictor and the LRS3 dataset for training.
The results are illustrated in Table 7. To assess the effectiveness of our proposed approach, we compared it with a previously introduced LLM-based VSR framework (VSP-LLM) Yeo et al. (2024a). VSP-LLM uses unit-based deduplication to reduce sequence length by up to 53%, reporting a WER of 26.7%. Although our method yields a slightly higher WER of 28.7%, it’s worth noting that it reduces the visual speech features per second by up to 23%.
5.4 Additional Results at Different SNR Levels
To facilitate fair comparison and follow-up research, we provide additional results for our proposed method at various SNR levels. The experiments were conducted at SNR levels of -15, -10, -5, 0, 5, 10, and 15 dB, following previous works. Table 8 illustrates the WERs at each SNR level.
Across the different SNR levels, the proposed method outperforms the previous approach Cheng et al. (2023a). Specifically, performance at 10dB to 15dB is comparable to that observed in a clean environment, while substantial improvements are evident between -10dB and 0dB.